An integer between 5 and 40 (both exclu Suppose sive) is chosen at random. What is the probability that 

i. It is an odd number?

ii. It is an even number and divisible by 4.


1
Expert's answer
2021-06-07T12:42:50-0400

i. In total, we have 34 integers from 5 to 40, so n=5.

The number of odd numbers from 5 to 40 is m=17

Then the wanted probability is "p = \frac{m}{n} = \frac{{17}}{{34}} = \frac{1}{2}"

Answer: "p = \frac{1}{2}"

ii. The number of even numbers divisible by 4 in the range from 5 to 40 is m=8, so

"p = \frac{m}{n} = \frac{{8}}{{34}} = \frac{4}{17}"

Answer: "p =\frac{4}{17}"